OpenStudy (anonymous):

Simplify tan(sin^−1 x)

OpenStudy (anonymous):

make a triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

it is just for a reference that is a picture of a triangle with an angle whose sine is \(x\)\you need the other side of the triangle, which you find by memory or pythagoaras it is \[\sqrt{1-x^2}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

tangent is opposite over adjacent

OpenStudy (anonymous):

i.e. \[\tan(\sin^{-1}(x))=\frac{x}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

hope it is more or less clear this gimmick works for all problems like \[\sin(\cos^{-1}(x))\] etc

OpenStudy (anonymous):

Thank You!